OpenStudy (anonymous):

what rate of interest compounded annually is required to double an investment in 3 years? please help!!!

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf A=P(1+\frac{r}{n} )^{nt} }\) P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year \(\Huge\color{blue}{ \sf 100=50(1+\frac{r}{1} )^{1\times 3 }}\) \(\Huge\color{blue}{ \sf 100=50(1+\frac{r}{1} )^{3 }}\) \(\Huge\color{blue}{ \sf 100=50(\frac{1+r}{1} )^{3 }}\) (in this step I divide both sides by 50 ) \(\Huge\color{blue}{ \sf 2=(\frac{1+r}{1} )^{3 }}\) \(\Huge\color{blue}{ \sf 2=\frac{(1+r)^3}{1^3} }\) \(\Huge\color{blue}{ \sf 2=\frac{(1+r)^3}{1} }\) \(\Huge\color{blue}{ \sf 2= (1+r)^3 }\) \(\Huge\color{blue}{ \sf \sqrt[3]{2^3} =(1+r)^3 }\) (in this step I take cube root of both sides ) ....

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf 2= (1+r)^3 }\) \(\Huge\color{blue}{ \sf \sqrt[3]{2} =(1+r) }\) (in this step I take cube root of both sides ) \(\Huge\color{blue}{ \sf \sqrt[3]{2} =1+r }\) \(\Huge\color{blue}{ \sf \sqrt[3]{2} -1=r }\) \(\Huge\color{blue}{ \sf (~~~~~ \sqrt[3]{2} ≈1.26 ~~~) }\)